Let f(x) = x2 + 3x – 4 and g(x) = x + 5. Find f(x) • g(x).
Answer by oscargut(891)

Solution:
f(x)g(x)=(x^2+3x-4)(x+5) = x^3+8x^2+11x-20
Answer: x^3+8x^2+11x-20

If you have any doubt or if you have more problems my e-mail is:
mthman@gmail.com